OpenStudy (samigupta8):

a car of mass M is moving horizontally on a circular path.At an instant its speed is v and is increasing at a rate a A.The acceleration of the car is toward the centre of the path B.The magnitude of the frictional force on the car is greater than mv^2/r C.The friction coefficient between the ground and the car is not less than a/g

OpenStudy (anonymous):

do you have any thoughts?!

OpenStudy (anonymous):

what is your question.s yes in a circular path acceleration is always going to point in towards the center.

OpenStudy (anonymous):

@a-higbee you may have overlooked this, this is not a uniform circular motion.. i think she wants to know which of them is right

OpenStudy (anonymous):

oh ok. sorry. let me re-look at the problem

OpenStudy (anonymous):

this link should answer your question. http://www.physicsclassroom.com/Class/circles/U6L1c.cfm

OpenStudy (samigupta8):

yes of couse i want to know of the three whcih one is correct and make it a note dat it's a multi correct answer so pls...accordingly give the answer with explanation

OpenStudy (anonymous):

I think the answer is B and C.  Thus \[at = \sqrt{(ar)^2 + (al)^2}\]. \[Frictional Force = \mu* (m*g)=m*(at)\] thus \[\mu = (at)/g\] \[(at > (al= a)) \therefore \mu > a/g\], also net force = frictional force(thats how vehicles move), net force = \[m*(at)\]

OpenStudy (samigupta8):

what is at here

OpenStudy (anonymous):

mentioned in the figure, its total acceleration...the vector addition of ar (angular acceleration ) and al( linear acceleration)..